OpenStudy (anonymous):

find the distance between points M(6,16) and Z(1,14) to the nearest tenth.

OpenStudy (aaronq):

this is math.. use distance formula .. |D| =\[\sqrt{(x1-x2)^2 + (y1-y2)^2}\]

OpenStudy (anonymous):

i dont understand

OpenStudy (aaronq):

you're finding the difference between the 2 points.. \[|MZ|= \sqrt{(6-1)^2+(16-14)^2} = \sqrt{29}\]

OpenStudy (anonymous):

so 29 is the answer?

OpenStudy (aaronq):

\[\sqrt{29} is the answer\]

OpenStudy (anonymous):

okay thanks.